### v2

# Graphical Scattering Equations

Abstract
The CHY scattering equations on the moduli space $\mathcal{M}_{0,n}$ play a prominent role at the interface of particle physics and algebraic statistics. We study the scattering correspondence when the Mandelstam invariants are restricted to a fixed graph on $n$ vertices.
1 Introduction
The CHY scattering potential in particle physics [2] is the following function in $n$ unknowns:
$L(x)\,\,\,=\,\,\sum_{1\leq i<j\leq n}\!\!s_{ij}\,{\rm log}(x_{i}-x_{j}).$ (1)
The coefficients $s_{ij}$ are parameters, known as Mandelstam invariants. They are assumed to satisfy $n$ momentum conservation relations. Setting $s_{ii}=0$ and $s_{ij}=s_{ji}$, these relations are
$\quad\sum_{j=1}^{n}s_{ij}\,=\,0\qquad\quad{\rm for}\,\,i=1,2,\ldots,n.$ (2)
The critical points of the potential $L(x)$ are the solutions to the CHY scattering equations
$\qquad\frac{\partial L}{\partial x_{i}}\,\,=\,\,\sum_{j=1}^{n}\frac{s_{ij}}{x_{i}-x_{j}}\,\,\,=\,\,\,0\qquad\quad{\rm for}\,\,i=1,2,\ldots,n.$ (3)
The hypothesis (2) ensures that (1) is invariant under linear fractional transformations
$x_{i}\,\,\mapsto\,\,\frac{a\,x_{i}+b}{c\,x_{i}+d}\qquad{\rm for}\,\,\,i=1,2,\ldots,n.$ (4)
Here $x_{i}$ is a point in $\mathbb{P}^{1}=\mathbb{C}\cup\{\infty\}$. The solution set to (3) consists of (n3)!(n-3)! orbits under the ${\rm PGL}(2)$ action in (4). Each orbit is a point in the moduli space $\mathcal{M}_{0,n}\subseteq(\mathbb{P}^{1})^{n}/\!/{\rm PGL}(2)$. Hence, the scattering potential $L(x)$ in (1) has (n3)!(n-3)! critical points in $\mathcal{M}_{0,n}$; see [2, 9, 15]. Each of them corresponds to a surface of degree n2n-2 in $\mathbb{P}^{n-1}$, given by an ideal as in (14).
In this paper we ask the following question: what happens when some $s_{ij}$ are set to zero? Will the solutions to (2) and (3) still be reasonable? This is the case in the study of minimal kinematics in [4]. We here offer a vast generalization. Let us begin with two small scenarios.
Example 1.1 ($n=4$).
Assuming $x_{i}\not=x_{j}$, the equations (2) and (3) are equivalent to
$\begin{matrix}&s_{12}-s_{34}\,\,=\,\,s_{13}-s_{24}\,\,=\,\,s_{14}-s_{23}\,\,=\,\,s_{12}+s_{13}+s_{14}&=&0\\ {\rm and}\quad&s_{34}x_{1}x_{2}+s_{24}x_{1}x_{3}+s_{14}x_{2}x_{3}+s_{23}x_{1}x_{4}+s_{13}x_{2}x_{4}+s_{12}x_{3}x_{4}&=&\,0.\end{matrix}\vskip-5.69046pt$
The quadric defines a surface in $\mathbb{P}^{3}$. It encodes a ${\rm PGL}(2)$ orbit in $(\mathbb{P}^{1})^{4}$, which is one point in $\mathcal{M}_{0,4}$. If we set one parameter $s_{ij}$ to zero, the system collapses, with no solutions in $\mathcal{M}_{0,4}$.
Example 1.2 ($n=5$).
The scattering potential has two critical points on the surface $\mathcal{M}_{0,5}$. There is one critical point if $s_{45}=0$, and also if $s_{23}=s_{45}=0$. On the other hand, the system collapses for $s_{34}=s_{45}=0$. There are no solutions when three $s_{ij}$ are zero. See Example 4.9.
To address our question, we fix a simple graph $G\subseteq\binom{[n]}{2}$ with vertices in $n]=\{1,\ldots,n\}$, and we set $s_{ij}=0$ for all non-edges $ij\not\in G$. Then (2) is a system of $n$ linear equations in the variables $s_{ij}$ for $ij\in G$. Its solution space in $\mathbb{P}^{|G|-1}$ is the kinematic space, denoted $\mathcal{K}_{G}$. The equations (3) are linear in $s_{ij}$, and they define a vector bundle over an open subset of $\mathbb{P}^{n-1}$. Its total space $\mathcal{V}_{G}$ is called the scattering correspondence of $G$. Its Zariski closure is an irreducible variety $\overline{\mathcal{V}}_{G}$ in $\mathcal{K}_{G}\times\mathbb{P}^{n-1}$. We want the graph $G$ to have enough edges so that the scattering correspondence $\mathcal{V}_{G}$ is non-empty. Moreover, if we consider $\mathcal{V}_{G}$ modulo the ${\rm PGL}(2)$ action (4), the two natural projections should have desirable properties:
$\mathcal{K}_{G}\,\longleftarrow\,\mathcal{V}_{G}\,\longrightarrow\,\mathcal{M}_{0,n}.$ (5)
For the complete graph $G=K_{n}$, we have ${\rm dim}(\mathcal{V}_{K_{n}})=2n-3$, the left map is (n3)!(n-3)!-to-one, and fibers on the right are linear subspaces of codimension n3n-3 in the kinematic space $\mathcal{K}_{G}$.
The graph $G$ is copious if it has enough edges to yield the properties above. By Examples 1.2 and 4.9, only three graphs with $n=5$ vertices are copious. We shall introduce four different approaches to what copious means. Our main theorem states that these four agree.
Theorem 1.3.
For every graph $G$, the following three conditions are equivalent:
  1. 1.
    $G$ is geometrically copious, (Definition 3.2)
  2. 2.
    $G$ is matroidally copious, (Definition 3.7)
  3. 3.
    $G$ is algebraically copious. (Definition 4.2)
These are also equivalent to the following condition for all graphs $G$ with a universal vertex:
  1. 4.
    $G$ is topologically copious. (Definition 5.2)
The implication from 4 to 1,2,3 holds for every graph $G$.
Developing these different notions of copious is our main thread. One important feature of the present paper is that all $n$ variables $x_{1},x_{2},\ldots,x_{n}$ are treated equally. We do not break their symmetry. In most other studies with explicit computations on $\mathcal{M}_{0,n}$, three of the points are fixed at $0,1,\infty$. That choice, called gauge fixing, is incompatible with our graph $G$, unless $G$ has a universal vertex (i.e., a vertex that is connected to all other vertices). For this reason, our commutative algebra takes place over a polynomial ring in $x_{1},x_{2},\ldots,x_{n}$.
We now discuss the organization of this article and its other main results. In Section 2 we examine the implications of momentum conservation. Proposition 2.2 rules out all graphs $G$ whose even-cycle matroid has a co-loop. Lemma 2.4 leads to an algebraic formulation of our problem which respects the ${\rm PGL}(2)$ action on the $n$ variables in (4). In Section 3 we take a closer look at the scattering correspondence $\mathcal{V}_{G}$. Its vector bundle structure defines the scattering matroid $\mathcal{S}(G)$, which is our object of study in Lemma 3.4 and Proposition 3.6. The section concludes with the proof that geometrically copious equals matroidally copious.
In Section 4 we turn to commutative algebra. Building on recent advances in [7, 10], we develop tools for computing the prime ideal $I_{G}$ of the scattering correspondence. This involves the equivalent concept of algebraically copious (Lemma 4.5). We introduce the multidegree of the ideal $I_{G}$, and show in Proposition 4.8 how this reveals the maximum likelihood degree $\mu(G)$. Proposition 4.11 offers a census of copious graphs with $n=6$ vertices.
In Section 5 we define a very affine variety $\mathcal{M}_{G}$ which is a partial compactification of the moduli space $\mathcal{M}_{0,n}$ associated with the graph $G$. This parallels the tropical construction due to Fry [5] on the moduli space of graphically stable rational curves. The signed Euler characteristic of $\mathcal{M}_{G}$ is the ML degree $\mu(G)$. This rests on Huh’s theorem in [6] and is the content of Theorem 5.3. Section 5 concludes with completing the proof of Theorem 1.3.
Section 6 is our contribution to experimental mathematics. We present exhaustive computations for graphs up to $n=9$ vertices, using both symbolic and numerical methods. Theorem 6.1 summarizes the findings. Our database of copious graphs is made available on Zenodo [1]. We illustrate how the data can be used to gain insights and generate conjectures.
Section 7 addresses the combinatorics of the ML degree. Conjecture 7.1 suggests a general formula. This is compared with the standard approach, using bounded regions and chromatic polynomials, which applies for special graphs that allow gauge fixing (Theorem 7.7).
Section 8 is devoted to two further topics. First, we generalize the logarithmic discriminant of [8, Section 8] from $K_{n}$ to arbitrary graphs $G$. Computational results on its degree $\delta(G)$ are given in Proposition 8.2. Second, we study graphs $G$ arising from hypertrees [3, 4]. We examine whether such $G$ are copious, and we explore scattering equations in the setting of [4, Section 5]. Our favorite hypertree is the bipyramid (Proposition 8.4 and Corollary 8.6). Ample discussions guide our readers to a plethora of fruitful directions for future research.
2 Momentum Conservation
Given a graph $G\subseteq\binom{[n]}{2}$, its kinematic space $\mathcal{K}_{G}$ is the linear subspace of $\mathbb{C}^{|G|}$ defined by
$\sum_{j\in G_{i}}s_{ij}\,=\,0\qquad{\rm for}\,\,i=1,2,\ldots,n.$ (6)
Here $G_{i}$ is the set of vertices of $G$ that are incident to the vertex $i$. As in the Introduction, we usually identify the kinematic space $\mathcal{K}_{G}$ with its image in the projective space $\mathbb{P}^{|G|-1}$.
The $n$ linear forms in (6) determine a matroid $M(\mathcal{K}_{G})$ of rank at most $n$ on the ground set $G$. Namely, $M(\mathcal{K}_{G})$ is the matroid of the vector configuration $\bigl\{e_{i}+e_{j}\in\mathbb{R}^{n}:ij\in G\bigr\}$. This matroid is known as the even cycle matroid of the graph $G$. Zaslavsky’s classical theory [16] refers to $M(\mathcal{K}_{G})$ as the signed graphic matroid of $G$ for negatively signed edges.
The following result is well-known at the interface of graph theory and matroid theory.
Lemma 2.1.
The circuits of the even cycle matroid $M(\mathcal{K}_{G})$ are
  1. 1.
    even cycles;
  2. 2.
    pairs of odd cycles intersecting in a single vertex;
  3. 3.
    pairs of vertex-disjoint odd cycles connected by a simple path.
We say that $G$ is compatible with momentum conservation if $\mathcal{K}_{G}\cap(\mathbb{C}^{*})^{|G|}$ is non-empty. This means that momentum conservation can be achieved with $s_{ij}$ that are supported on $G$.
Proposition 2.2.
A graph $G$ is compatible with momentum conservation if and only if its even cycle matroid $M(\mathcal{K}_{G})$ has no co-loop. This holds if and only if $G$ has no degree one vertices and no edges that belong to every odd cycle in some connected component of G\,G.
Proof.
The graph $G$ is not compatible with momentum conservation if and only if $\mathcal{K}_{G}$ lies in some coordinate hyperplane $\{s_{ij}=0\}$ of $\mathbb{C}^{|G|}$. This means that $ij$ is a co-loop of $M(\mathcal{K}_{G})$. The second statement can be derived from Lemma 2.1. In Zaslavsky’s framework, it follows from the fact that the balanced subgraphs are the bipartite subgraphs [16, Corollary 7.D.1]. ∎
Corollary 2.3.
The codimension of $\mathcal{K}_{G}$ equals the rank of $M(\mathcal{K}_{G})$. This equals $n$ minus the number of bipartite connected components of $G$, where isolated vertices are viewed as bipartite.
We shall use the following abbreviation for the codimension of the kinematic space:
$\kappa(G)\,\,:=\,\,{\rm codim}(\mathcal{K}_{G})\,\,=\,\,{\rm rank}(M(\mathcal{K}_{G})).$ (7)
Our object of study in this paper is the graphical scattering potential
$L_{G}(x)\,\,=\,\,\sum_{ij\in G}s_{ij}\,{\rm log}(x_{i}-x_{j}).$
The Mandelstam parameters $s_{ij}$ are non-zero complex numbers. We assume that (6) holds. For this to make sense, the graph $G$ must be compatible with momentum conservation.
Lemma 2.4.
The scattering potential $L_{G}$ is constant on the orbits of the ${\rm PGL}(2)$ action (4). If we fix a branch of the complex logarithm, then $L_{G}$ is a well-defined function on $\mathcal{M}_{0,n}$.
Proof.
We consider the lattice $\mathcal{K}_{G}\,\cap\,\mathbb{Z}^{|G|}$ of all integer vectors in the kinematic space. This is a free abelian group of rank $r:=|G|-\kappa(G)$. Every vector $u\in\mathcal{K}_{G}\,\cap\,\mathbb{Z}^{|G|}$ gives rise to a character $\tilde{u}:(\mathbb{C}^{*})^{|G|}\rightarrow\mathbb{C}^{*}$. We write $\tilde{u}$ as a Laurent monomial in the quantities $x_{i}-x_{j}$ for $ij\in G$. Then $\tilde{u}$ is a rational function on $\mathbb{P}^{n-1}$ which is constant on ${\rm PGL}(2)$ orbits. We call $\tilde{u}$ the generalized cross-ratio, and we view it as a regular function on the moduli space $\mathcal{M}_{0,n}$.
Let $u_{1},u_{2},\ldots,u_{r}$ be a basis of $\mathcal{K}_{G}\,\cap\,\mathbb{Z}^{|G|}$. Since $s=(s_{ij})$ satisfies momentum conservation, we can write $s=t_{1}u_{1}+t_{2}u_{2}+\cdots+t_{r}u_{r}$, where the $t_{k}$ are coordinates on $\mathcal{K}_{G}$. We then have
$L_{G}(x)\,\,=\,\,\sum_{k=1}^{r}t_{k}\cdot{\rm log}(\tilde{u}_{k}(x)).$ (8)
This is an expression in terms of generalized cross-ratios, so it is ${\rm PGL}(2)$ invariant.∎
The following graph appears prominently in [4, Section 5] and in [7, Theorem 5.2].
Example 2.5 (Octahedron).
Let $G$ be the graph with $n=6$ and non-edges $12,34,56$. Here $r=6$, and we choose the following integer basis for the kinematic subspace $\mathcal{K}_{G}$ in $\mathbb{R}^{|G|}=\mathbb{R}^{12}$:
$\begin{matrix}u_{1}\,=\,e_{13}+e_{24}-e_{14}-e_{23},&u_{2}\,=\,e_{15}+e_{26}-e_{16}-e_{25},&u_{3}\,=\,e_{35}+e_{46}-e_{36}-e_{45},\\ u_{4}\,=\,e_{14}+e_{25}-e_{15}-e_{24},&u_{5}\,=\,e_{23}+e_{45}-e_{24}-e_{35},&u_{6}\,=\,e_{25}+e_{46}-e_{26}-e_{45}.\end{matrix}$ (9)
We can thus write $s=\sum_{k=1}^{6}t_{k}u_{k}$ for $s\in\mathcal{K}_{G}$. The corresponding cross-ratios are
$\begin{matrix}\tilde{u}_{1}(x)\,=\,\frac{(x_{1}-x_{3})(x_{2}-x_{4})}{(x_{1}-x_{4})(x_{2}-x_{3})},&\tilde{u}_{2}(x)\,=\,\frac{(x_{1}-x_{5})(x_{2}-x_{6})}{(x_{1}-x_{6})(x_{2}-x_{5})},&\tilde{u}_{3}(x)\,=\,\frac{(x_{3}-x_{5})(x_{4}-x_{6})}{(x_{3}-x_{6})(x_{4}-x_{5})},\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \tilde{u}_{4}(x)\,=\,\frac{(x_{1}-x_{4})(x_{2}-x_{5})}{(x_{1}-x_{5})(x_{2}-x_{4})},&\tilde{u}_{5}(x)\,=\,\frac{(x_{2}-x_{3})(x_{4}-x_{5})}{(x_{2}-x_{4})(x_{3}-x_{5})},&\tilde{u}_{6}(x)\,=\,\frac{(x_{2}-x_{5})(x_{4}-x_{6})}{(x_{2}-x_{6})(x_{4}-x_{5})}.\end{matrix}$ (10)
In terms of these six cross-ratios, the scattering potential on the octahedron graph $G$ equals
$L_{G}(x)\,=\,\sum_{ij\in G}s_{ij}\,{\rm log}(x_{i}-x_{j})\,=\,\sum_{k=1}^{6}t_{k}\,{\rm log}(\tilde{u}_{k}(x)).$ (11)
This demonstrates how $L_{G}(x)$ can be written manifestly as a function on $\mathcal{M}_{0,6}$, but without breaking the symmetry among the $n$ points. As was said before, we abstain from gauge fixing.
For any graph $G$ on $n$ vertices, we wish to solve the graphical scattering equations
$\frac{\partial{L_{G}}}{\partial x_{1}}\,=\,\frac{\partial{L_{G}}}{\partial x_{2}}\,=\,\,\cdots\,\,=\,\frac{\partial{L_{G}}}{\partial x_{n}}\,=\,0.$ (12)
[…]
$x_{i}\,=\,\frac{au_{i}+b}{cu_{i}+d},\qquad{\rm with}\,\,\,cu_{i}+d\neq 0,\qquad{\rm for}\,\,\,i=1,2,\ldots,n.$ (13)
By eliminating $a,b,c,d$ from the $n$ equations in (13), we obtain the determinantal ideal
$I_{4}\!\small\begin{pmatrix}1&1&\cdots&1\\ u_{1}&u_{2}&\cdots&u_{n}\\ x_{1}&x_{2}&\cdots&x_{n}\\ u_{1}x_{1}\!&\!u_{2}x_{2}&\!\cdots&\!u_{n}x_{n}\end{pmatrix}\,=\,\,I_{2}\!\small\begin{pmatrix}[13]\langle 23\rangle&[14]\langle 24\rangle&\cdots&[1n]\langle 2n\rangle\vskip 3.0pt plus 1.0pt minus 1.0pt\\ [23]\langle 13\rangle&[24]\langle 14\rangle&\cdots&[2n]\langle 1n\rangle\\ \end{pmatrix}\,\subset\,\mathbb{C}[x_{1},\ldots,x_{n}].$ (14)
The entries of the $2\times(n-2)$ matrix on the right are written in physics-inspired notation:
$ij]=x_{i}-x_{j}\qquad{\rm and}\qquad\langle ij\rangle=u_{i}-u_{j}.$
The homogeneous ideal in (14) is prime. That ideal is our algebraic encoding for one point $u$ in the moduli space $\mathcal{M}_{0,n}$. The variety defined by (14) in $\mathbb{P}^{n-1}$ is a surface of degree n2n-2. This surface is the projectivized affine chart of the orbit closure of the point $(u_{1},\ldots,u_{n})\in(\mathbb{P}^{1})^{n}$.
We expect the scattering equations (12) to have finitely many complex solutions in $\mathcal{M}_{0,n}$, for fixed $s\in\mathcal{K}_{G}$. This expectation will be made precise in the next section. The number of solutions is the ML degree $\mu(G)$. Thus, the variety defined by (12) in $\mathbb{P}^{n-1}$ consists of $\mu(G)$ surfaces of degree n2n-2, each given by a prime ideal (14). For instance, for the octahedron $G$ from Example 2.5, we find exactly $\mu(G)=2$ such surfaces in $\mathbb{P}^{5}$. See [4, Example 5.4].
We close this section by highlighting the similarities and differences between [7] and the present paper. Let $\mathcal{H}_{G}$ be the graphical arrangement, obtained from $\mathcal{H}$ by deleting all hyperplanes with label $ij\notin G$. The likelihood correspondence of the arrangement $\mathcal{H}_{G}$ is studied in [7, Section 4]. The similarity is that both theories concern the same function $L_{G}$.
The crucial difference lies in the assumptions on the parameters $s_{ij}$. In [7] one has $\sum_{1\leq i<j\leq n}s_{ij}=0$, but otherwise the $s_{ij}$ are unconstrained. This results in finitely many critical points in $\mathbb{P}^{n-1}$. By contrast, we here assume momentum conservation (6). This assumption implies that the critical points in $\mathbb{P}^{n-1}$ are not isolated, but they come in ${\rm PGL}(2)$ orbits. These are the surfaces defined by (14), representing points in the moduli space $\mathcal{M}_{0,n}$.
The likelihood ideal in [7] is usually generated by Mühlherr’s equations from [10]. We refer to [7, Theorem 5.2], which says that the octahedron is the only exception for $n=6$. Mühlherr’s theorem in [10] is relevant here, and will be discussed in Section 4. However, the commutative algebra of our scattering ideals is certainly more complicated than that in [7].
3 The Scattering Matroid
[…]
$\mathcal{V}_{G}\,\,=\,\,\left\{(s,x)\in\mathbb{P}^{|G|-1}\times\mathbb{P}^{n-1}\backslash\mathcal{H}:\,\frac{\partial L_{G}}{\partial x_{i}}=0\,\,\,{\rm and}\,\,\sum\limits_{j\in G_{i}}s_{ij}=0\ \text{ for }\ i=1,\dots,n\right\}.$ (15)
We write $\overline{\mathcal{V}}_{G}$ for the Zariski closure of $\mathcal{V}_{G}$ in $\mathcal{K}_{G}\times\mathbb{P}^{n-1}$. We frequently pass from $\mathbb{P}^{n-1}\backslash\mathcal{H}$ to $\mathcal{M}_{0,n}$ by taking the quotient modulo ${\rm PGL}(2)$. We thus view $\mathcal{V}_{G}$ as a subvariety of $\mathcal{K}_{G}\times\mathcal{M}_{0,n}$.
The scattering correspondence $\mathcal{V}_{G}$ is defined by rational function equations (12). It is far from obvious how these can be replaced with polynomials. How can we clear denominators gracefully? This will be addressed in Section 4. For a first glimpse of that issue, consider the following polynomials of bidegree $(1,k)$ in $(s,x)$. They are symmetric functions in $(x_{i},x_{j})$:
$\qquad\sum\limits_{ij\in G}\left(s_{ij}\sum\limits_{\ell=0}^{k}x_{i}^{k-\ell}x_{j}^{\ell}\right),\qquad{\rm where}\,\,k=2,3,\ldots,n-2.$ (16)
For $k=1$, this is a linear form in $x_{1},\ldots,x_{n}$ whose coefficients are precisely the $n$ momentum conservation relations (6). Hence, it vanishes on $\mathcal{V}_{G}$. The same is true for $k\geq 2$ as well.
Lemma 3.1.
The polynomials in (16) vanish on the scattering correspondence $\mathcal{V}_{G}$.
Proof.
Up to an index shift, these are the ideal generators $\theta_{k}$ in [7, Theorem 5.8]. ∎
We now come to the first definition of copious, referring to a graph with enough edges.
Definition 3.2.
A graph $G\subseteq\binom{[n]}{2}$ is geometrically copious if the projection of the scattering correspondence $\mathcal{V}_{G}$ to $\mathcal{M}_{0,n}$ is dominant and the generic fiber over $\mathcal{K}_{G}$ has dimension two. In the formulation of (5), the latter condition means that the map on the left is finite-to-one.
If $G$ is geometrically copious then its ML degree $\mu(G)$ is a well-defined positive integer. It is the degree of the map on the left in (5), i.e., the number of critical points of $L_{G}(x)$ in $\mathcal{M}_{0,n}$.
Proposition 3.3.
Every vertex in a geometrically copious graph $G$ has degree at least three.
Proof.
Consider a vertex $i$ of degree $\leq 2$ in $G$. If $G_{i}=\{j,\ell\}$ then (6) implies $s_{ij}=-s_{i\ell}$, and ${\partial L_{G}}/{\partial x_{i}}=0\,$ forces $x_{j}=x_{\ell}$. Hence the map $\mathcal{V}_{G}\rightarrow\mathcal{M}_{0,n}$ is not dominant. If $G_{i}\subseteq\{\ell\}$ then the variable $x_{i}$ does not occur in the scattering equations, so the fibers of $\mathcal{K}_{G}\leftarrow\mathcal{V}_{G}$ are either empty or positive-dimensional. In both cases, $G$ is not geometrically copious. ∎
A key role in what makes a graph $G$ copious is played by matroids on the ground set $G$. We saw a first instance of this in Proposition 2.2. These matroids furnish the combinatorial framework for $\mathcal{V}_{G}$ as a vector bundle over $\mathcal{M}_{0,n}$. The point is that the equations (6) are linear over $\mathbb{Z}$, and the equations (12) are linear over the field $\mathbb{C}(x)$. We define the scattering matroid $\mathcal{S}(G)$ to be the matroid on $G$ that is given by the $2n$ linear equations in (6) and (12).
Lemma 3.4.
The scattering matroid $\mathcal{S}(G)$ has rank at most 2n32n-3. This upper bound is attained for the complete graph $K_{n}$. In symbols, we have ${\rm rank}(\mathcal{S}(K_{n}))=2n-3$.
Proof.
The fiber in $\mathcal{V}_{G}$ over a generic point $u$ in $\mathbb{P}^{n-1}$ is the projectivization of a subspace of the vector space $\mathbb{C}^{|G|}$. By definition, $\mathcal{S}(G)$ is the matroid defined by this subspace. We represent $\mathcal{S}(G)$ by the following $2n\times|G|$ matrix $S_{G}$ with entries in the field $\mathbb{C}(x_{1},\dots,x_{n})$. The columns of $S_{G}$ are indexed by edges $ij\in G$. The rows of $S_{G}$ are the coefficients of (6) and (12) with respect to the $s_{ij}$. Explicitly, the entries of $S_{G}$ in a given column $ij$ are as follows. For $k=1,\ldots,n$, the entry in row $k$ is $1$ if $k\in\{i,j\}$ and it is 0 otherwise. The entry in row $n+k$ is $(x_{i}-x_{j})^{-1}$ if $i=k$, the entry is $(x_{j}-x_{i})^{-1}$ if $j=k$, and it is 0 otherwise.
Fix the complete graph $G=K_{n}$, and consider the $2n\times\binom{n}{2}$ matrix $S_{K_{n}}$ defined above. Since $I_{K_{n}}$ is a complete intersection by [7, Example 2.1], generated by (6) and the n3n-3 polynomials in (16), we can replace the last $n$ rows of our matrix $S_{K_{n}}$ by the n3n-3 coefficient vectors of (16). The resulting $(2n-3)\times\binom{n}{2}$ matrix has linearly independent rows, and it has the same row space as $S_{K_{n}}$. It therefore defines the same matroid $\mathcal{S}(K_{n})$ of rank 2n32n-3.
The matrix $S_{G}$ for any subgraph $G$ of $K_{n}$ is obtained from the $(2n-3)\times\binom{n}{2}$ matrix $S_{K_{n}}$ by deleting all columns corresponding to non-edges of $G$. The rank of a matrix cannot increase if we delete some columns. Hence $S_{G}$ has rank at most 2n32n-3, and so does $\mathcal{S}(G)$. ∎
The proof suggests that we start by examining the matrix $S_{K_{n}}$ for the complete graph. For any graph $G$, the matrix $S_{G}$ is obtained by deleting the columns of $S_{K_{n}}$ indexed by non-edges of $G$. The codimension $\kappa(G)$ in (7) is the rank of the submatrix of $S_{G}$ given by the first $n$ rows. For $G$ to be copious, the whole matrix $S_{G}$ should have rank $\kappa(G)+n-3$.
Example 3.5 ($n=6$).
The matrix $S_{K_{6}}$ has $12$ rows and $15$ columns, and it equals
$\footnotesize\begin{bmatrix}1&1&1&1&1&0&\cdots&0&0\\ 1&0&0&0&0&1&\cdots&0&0\\ 0&1&0&0&0&1&\cdots&0&0\\ 0&0&1&0&0&0&\cdots&1&0\\ 0&0&0&1&0&0&\cdots&0&1\\ 0&0&0&0&1&0&\cdots&1&1\\ (x_{1}{-}x_{2})^{{-}1}\!&\!(x_{1}{-}x_{3})^{{-}1}\!&\!\!(x_{1}{-}x_{4})^{{-}1}\!&\!(x_{1}{-}x_{5})^{{-}1}\!&\!\!(x_{1}{-}x_{6})^{{-}1}\!&0&\cdots&0&0\\ (x_{2}{-}x_{1})^{{-}1}\!&0&0&0&0&\!\!\!(x_{2}{-}x_{3})^{{-}1}&\cdots&0&0\\ 0&\!(x_{3}{-}x_{1})^{{-}1}\!&0&0&0&\!\!\!(x_{3}{-}x_{2})^{{-}1}&\cdots&0&0\\ 0&0&\!\!(x_{4}{-}x_{1})^{{-}1}\!&0&0&0&\cdots&(x_{4}{-}x_{6})^{{-}1}&0\\ 0&0&0&\!(x_{5}{-}x_{1})^{{-}1}\!&0&0&\cdots&0&\!\!\!\!(x_{5}{-}x_{6})^{{-}1}\\ 0&0&0&0&\!\!(x_{6}{-}x_{1})^{{-}1}\!&0&\cdots&0&\!\!\!\!(x_{6}{-}x_{5})^{-1}\\ \end{bmatrix}$.
Proposition 3.6.
If $G$ is geometrically copious then the following constraints hold:
[…]
${\rm and}\qquad\sum_{u\in\mathbb{P}^{n-1}\backslash\mathcal{H}}\!{\rm kernel}(S_{G}(u))\,\,=\,\,\mathcal{K}_{G}.$ (18)
The last condition is equivalent to
${\rm dim}\,\bigl(\bigcap_{u\in\mathbb{P}^{n-1}\backslash\mathcal{H}}{\rm rowspace}(S_{G}(u))\,\bigr)\,\,=\,\,\kappa(G).$ (19)
Proof.
We work in the setting of (5), where $\mathcal{V}_{G}\subset\mathcal{K}_{G}\times\mathcal{M}_{0,n}$. Our hypothesis says that the map to $\mathcal{K}_{G}$ is dominant and finite-to-one. This means that the fiber over a generic point in $\mathcal{M}_{0,n}$ does not lie in a proper subspace of $\mathcal{K}_{G}$. In particular, that fiber doesn’t lie in a coordinate subspace $\{s_{ij}=0\}$, that is, no edge $ij$ of $G$ is a co-loop of the scattering matroid $\mathcal{S}(G)$. In symbols, this is the first equality of (17). The fact that the map $\mathcal{V}_{G}\to\mathcal{K}_{G}$ is of finite degree also implies that the codimension of the scattering correspondence in $\mathcal{K}_{G}\times\mathcal{M}_{0,n}$ equals $n-3={\rm dim}(\mathcal{M}_{0,n})$. Its codimension in $\mathbb{P}^{|G|-1}\times\mathcal{M}_{0,n}$ is therefore ${\rm codim}(\mathcal{K}_{G})+n-3=\kappa(G)+n-3$. The latter codimension is precisely the rank of $\mathcal{S}(G)$, thus (17) holds.
[…]
Our second notion of what it means for a graph $G$ to be copious rests on matroid theory.
Definition 3.7.
A graph $G\!\subseteq\!\binom{[n]}{2}$ is called matroidally copious if both (17) and (18)=(19) are satisfied. In particular, the rank of the scattering matroid $\mathcal{S}(G)$ satisfies the upper bound
$\kappa(G)+n-3<|G|.$ (20)
Lemma 3.8.
A graph is matroidally copious if and only if it is geometrically copious.
Proof.
Proposition 3.6 states that conditions (17) and (18)=(19) are equivalent to the map $\mathcal{V}_{G}\to\mathcal{K}_{G}$ being finite-to-one. Indeed, (17)-(18) implies that the projection to $\mathcal{K}_{G}$ is dominant. In particular, generic fibers are non-empty. Since, by (17), we have $\dim\mathcal{V}_{G}=\dim\mathcal{K}_{G}$, this map is also of finite degree. To show that the map to $\mathcal{M}_{0,n}$ is dominant, the fiber over a generic point $u\in\mathbb{P}^{n-1}$ should be non-empty. This follows from (17). The corresponding linear system should have a nonzero solution. Equivalently, $\mathrm{rank}(S_{G}(u))<\min(|G|,2n)$ for a generic $u\in\mathbb{P}^{n-1}$. The left-hand side is the rank of the scattering matroid $\mathcal{S}(G)$, and by Lemma 3.4 this rank is bounded by 2n32n-3. Thus, it is enough to have $\mathrm{rank}(\mathcal{S}(G))<|G|$. ∎
[…]
4 Commutative Algebra
In this section we study the vanishing ideal $I_{G}$ of the scattering correspondence $\mathcal{V}_{G}$. We call this the scattering ideal. It lives in the polynomial ring $\mathbb{C}[s,x]$ with variables $s_{ij}$ for the edges in $G$ and variables $x_{1},\ldots,x_{n}$ for the vertices of $G$. This is the homogeneous coordinate ring of $\mathbb{P}^{|G|-1}\times\mathbb{P}^{n-1}$. The prime ideal $I_{G}$ is homogeneous in the $\mathbb{Z}^{2}$-grading on $\mathbb{C}[s,x]$ given by
${\rm deg}(s_{ij})=(1,0)\quad{\rm and}\quad{\rm deg}(x_{i})=(0,1).$ (21)
Let $K_{G}$ denote the ideal generated by the $n$ linear forms in (6) which express momentum conservation. As a first approximation to $I_{G}$, we write $I^{(0)}_{G}$ for the ideal generated by $K_{G}$ and
$\prod_{j\in G_{i}}(x_{i}-x_{j})\cdot\frac{\partial L_{G}}{\partial x_{i}}\qquad{\rm for}\,\,i\in[n].$ (22)
Proposition 4.1.
[…]
Proof.
The ideal $I_{G}$ is prime because it is the radical ideal of a vector bundle over a Zariski open subset of $\mathbb{P}^{n-1}$. In the definition of $\mathcal{V}_{G}$, the point $s$ is generic in $\mathcal{K}_{G}$ and $x$ is generic in $\mathbb{P}^{n-1}\backslash\mathcal{H}$. This implies that $I_{G}$ contains no non-zero polynomial in $x$, and the only polynomials in $s$ that appear in $I_{G}$ are those in $K_{G}$. From polynomials in the ideal $I^{(0)}_{G}$, we therefore remove any such extraneous factors in $x$ or $s$ alone. After eliminating $\kappa(G)$ of the Mandelstam variables $s_{ij}$ from the linear ideal $K_{G}$, we can write our requirements as follows:
$I_{G}\,\,=\,\,\bigl(\mathbb{C}(s)[x]\cdot I^{(0)}_{G}\,+\,\mathbb{C}(x)[s]\cdot I^{(0)}_{G}\bigr)\,\cap\,\mathbb{C}[s,x].$ (23)
[…]
Definition 4.2.
A graph $G\subseteq\binom{[n]}{2}$ is algebraically copious if $G$ is compatible with momentum conservation (6) and the ideal $I_{G}$ defined in (23) is a proper prime ideal in $\mathbb{C}[s,x]$.
Let $I^{(1)}_{G}$ be the ideal generated by $I^{(0)}_{G}$ and the polynomials in (16). This is a subideal of $I_{G}$. Proposition 4.1 holds verbatim for $I^{(1)}_{G}$. In some cases, we are now done, i.e. $I^{(1)}_{G}=I_{G}$.
Example 4.3.
The complete graph $K_{n}$ is algebraically copious. The ideal $I^{(0)}_{K_{n}}$ is generated by $n$ polynomials in (6) of degree $(1,0)$ and $n$ polynomials in (22) of degree (1,n2)(1,n\!-\!2). The ideal
$K_{K_{n}}+\langle\,x_{i}-x_{j}\,:\,1\leq i\!<\!j\leq n\rangle\,$
is an embedded prime of $I^{(0)}_{K_{n}}$. Saturating by this embedded prime, we arrive at $I_{G}^{(1)}=I_{G}$. Indeed, it follows from freeness of the braid arrangement [7, Example 2.1] that $I_{G}$ is the complete intersection ideal generated by $K_{K_{n}}$ and polynomials (16) for $k=2,3,\ldots,n-2$.
To illustrate Definition 4.2, we next show two graphs that are not algebraically copious.
Example 4.4 ($n=6$).
Let $G_{1}$ be the graph with non-edges $\{35,36,45,46\}$, and let $G_{2}$ be the graph with non-edges $\{16,24,25,34,35\}$. Using Macaulay2 computations, we find that
$\begin{matrix}&(x_{1}-x_{2})(x_{3}-x_{5})(x_{3}-x_{6})(x_{4}-x_{5})(x_{4}-x_{6})(s_{15}-s_{26})\in I_{G_{1}}^{(1)}&{\rm but}&s_{15}-s_{26}\not\in K_{G_{1}},\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \!{\rm and}\!\!&(x_{1}-x_{6})(x_{2}-x_{4})(x_{2}-x_{5})(x_{3}-x_{4})(x_{3}-x_{5})(s_{12}-s_{36})\in I_{G_{2}}^{(1)}&{\rm but}&s_{12}-s_{36}\not\in K_{G_{2}}.\end{matrix}\vskip-5.12128pt$
This shows that $K_{G_{i}}$ is strictly contained in the intersection of $\mathbb{C}(x)[s]\cdot I_{G_{i}}^{(1)}$ with $\mathbb{C}[s]$. The scattering equations have no solution in $\mathcal{M}_{0,6}$ for generic $s\in\mathcal{K}_{G_{i}}$, thus $\mathcal{V}_{G_{i}}=\emptyset$ for $i=1,2$.
Following Lemma 3.8, we now come to the next step in our proof of Theorem 1.3. The next lemma completes the equivalence of Definitions 3.2, 3.7 and 4.2. Hence, from now on, we simply use the adjective copious for a graph that satisfes any of these three definitions.
Lemma 4.5.
A graph $G$ is geometrically copious if and only if it is algebraically copious.
Proof.
We first proof the “only if” direction. The projection $\mathcal{V}_{G}\to\mathbb{P}^{n-1}$ is dominant if and only if the elimination ideal $I_{G}\cap\mathbb{C}[x]$ is zero. Hence, the corresponding conditions in the definitions of geometrically and algebraically copious graphs are equivalent. Now assume that the map $\mathcal{K}_{G}\leftarrow\mathcal{V}_{G}$ is finite-to-one. In particular, it is dominant onto $\mathcal{K}_{G}$, which is equivalent to $I_{G}\cap\mathbb{C}[s]=K_{G}$. Conversely, if a graph is algebraically copious, then $\mathcal{K}_{G}\leftarrow\mathcal{V}_{G}$ must also be finite-to-one. Indeed, if the map is not finite-to-one, there exists a variable, say $x_{i}$, which assumes infinitely many values in the scattering correspondence. We fix generic values $x_{2}=u_{2},\dots,x_{n}=u_{n}$ in $\mathbb{C}$, and we consider the matrix $S_{G}(x_{1},u_{2},\dots,u_{n})$ which represents the scattering matroid. For $x_{1}$ near $u_{2}$, the column $12$ has two arbitrarily large entries. Hence the kernel of the matrix lies in $\{s_{12}=0\}\subseteq\mathcal{K}_{G}$. This contradicts $I_{G}\cap\mathbb{C}[s]=K_{G}$. ∎
We next introduce one more ideal for approximating $I_{G}$. This is based on the connection with modules of derivations on graphic arrangements [7, Section 4]. The derivation module of $G$ is a $\mathbb{C}[x]$-module. It consists of all vectors $f(x)=(f_{1}(x),\ldots,f_{n}(x))\in\mathbb{C}[x]^{n}$ such that
$\qquad\,\sum_{i=1}^{n}f_{i}(x)\,\frac{\partial L_{G}}{\partial x_{i}}\qquad\hbox{is a polynomial in $\,\mathbb{C}[s,x]$}.$ (24)
[…]
$\qquad f_{i}(x)\,\,=\,\,\prod\limits_{t\in T}(x_{i}-x_{t})\,\,\,\,\text{ if }i\in C,\quad 0\ \text{otherwise}.$ (25)
Then (24) holds by [10, Lemma 3.4], and we obtain a polynomial of degree (1,|T|1)(1,|T|-1) in $I_{G}$.
Example 4.6.
[…]
We define $I^{(2)}_{G}$ to be the ideal in $\mathbb{C}[s,x]$ that is generated by $I^{(1)}_{G}$ and the polynomials (24) for all pairs $(C,T)$ as above. We summarize our discussion in the following corollary.
Corollary 4.7.
The ideal $I^{(2)}_{G}$ is contained in $I_{G}$. It also satisfies the conclusion of Proposition 4.1, namely: if $I^{(2)}_{G}$ is prime, $I^{(2)}_{G}\cap\mathbb{C}[x]=\{0\}$ and $I^{(2)}_{G}\cap\mathbb{C}[s]=K_{G}$, then $I^{(2)}_{G}=I_{G}$.
An important invariant of any $\mathbb{Z}^{2}$-homogeneous ideal in $\mathbb{C}[s,x]$ is its multidegree. For the scattering ideal $I_{G}$ this is the class of $\overline{\mathcal{V}}_{G}$ in the cohomology ring of $\mathbb{P}^{|G|-1}\times\mathbb{P}^{n-1}$. We have
$\qquad[\,\overline{\mathcal{V}}_{G}\,]\,\,=\,\,\sum_{i=0}^{c}\gamma_{i}\,\sigma^{i}\tau^{c-i}\,\,\,\,\in\,\,\mathbb{Z}[\sigma,\tau]/\langle\,\sigma^{|G|},\,\tau^{n}\,\rangle\,\,=\,\,H^{*}\bigl(\mathbb{P}^{|G|-1}\times\mathbb{P}^{n-1},\mathbb{Z}\bigr).$ (26)
Here $c={\rm codim}(\mathcal{V}_{G})$, and $\gamma_{i}=\#\bigl(\mathcal{V}_{G}\cap(L_{i}\!\times\!M_{c-i})\bigr)$, where $L_{i}$ is a general subspace of dimension $i$ in $\mathbb{P}^{|G|-1}$ and $M_{c-i}$ is a subspace of dimension cic-i in $\mathbb{P}^{n-1}$. See Lemma 4.10.
Proposition 4.8.
Let $G$ be a copious graph. Then the sum in (26) for $\overline{\mathcal{V}}_{G}]$ runs from $i=\kappa$ to $c=\kappa+n-3$, where $\kappa=\kappa(G)$ is the codimension of the kinematic space $\mathcal{K}_{G}$ in $\mathbb{P}^{|G|-1}$. The last coefficient is $\gamma_{n+\kappa-3}=1$, and the first nonzero coefficient gives the ML degree:
$\mu(G)\,\,=\,\,\frac{\gamma_{\kappa}}{n-2}.$ (27)
Proof.
The kinematic space $\mathcal{K}_{G}$ has dimension $|G|-\kappa-1$. For copious graphs, the general fiber of the map $\mathcal{K}_{G}\leftarrow\mathcal{V}_{G}$ has dimension $2$. This implies ${\rm dim}(\mathcal{V}_{G})=|G|-\kappa+1$, and thus $c=n+\kappa-3$. Moreover, that fiber is the union of $\mu(G)$ surfaces, each representing one critical point of $L_{G}$ in $\mathcal{M}_{0,n}$. This union is a surface of degree $\gamma_{\kappa}=\mu(G)\cdot(n-2)$. The rest of the statement is clear. Note that since $G$ is matroidally copious, we have $c<|G|$. ∎
For instance, using the complete intersection in Example 4.3, the complete graph $K_{n}$ has
$\overline{\mathcal{V}}_{K_{n}}]\,\,=\,\,\sigma^{n}\prod_{k=2}^{n-2}(\sigma+k\tau)\,\,=\,\,(n-2)!\,\sigma^{n}\tau^{n-3}\,+\,\,\cdots\,\,+\,\frac{n(n-3)}{2}\,\sigma^{2n-4}\tau\,+\,\sigma^{2n-3}.$
From equation (27) we now see the well-known formula $\mu(K_{n})\,\,=\,\,(n-2)!/(n-2)\,\,=\,\,(n-3)!$.
Example 4.9 ($n=5$).
For each copious graph $G$, we display the ideal $I_{G}$ and its multidegree:
  • (a)
    The complete graph $G=K_{5}$ has $\overline{\mathcal{V}}_{K_{5}}]=6\sigma^{5}\tau^{2}+5\sigma^{6}\tau+\sigma^{7}$, and its scattering correspondence $\mathcal{V}_{K_{5}}$ is the complete intersection given by (16) for $k=2$ and $k=3$.
  • (b)
    If $G$ has one non-edge $45$, then $\overline{\mathcal{V}}_{G}]=3\sigma^{5}\tau^{2}+4\sigma^{6}\tau+\sigma^{7}$. The ideal has three minimal generators, two of degree $(1,2)$ and one of degree $(2,2)$. We take (16) for $k=2$, and
    $\begin{matrix}&&(x_{1}-x_{2})(x_{3}-x_{4})s_{24}\,\,+\,\,(x_{1}-x_{3})(x_{2}-x_{4})s_{34},\\ &{\rm and}&(x_{2}-x_{4})(x_{3}-x_{5})s_{25}s_{34}\,-\,(x_{2}-x_{5})(x_{3}-x_{4})s_{24}s_{35}.\end{matrix}$
  • (c)
    If $G$ has two non-edges $23$ and $45$, then $\overline{\mathcal{V}}_{G}]=3\sigma^{5}\tau^{2}+3\sigma^{6}\tau+\sigma^{7}$. The ideal $I_{G}$ has three minimal generators modulo $K_{G}$, all of degree $(1,2)$. We take (16) for $k=2$, and
    $\begin{matrix}&&(x_{1}-x_{2})(x_{3}-x_{5})s_{25}\,+\,(x_{1}-x_{3})(x_{2}-x_{5})s_{35},\\ &{\rm and}&(x_{1}-x_{4})(x_{3}-x_{5})s_{34}\,+\,(x_{1}-x_{5})(x_{3}-x_{4})s_{35}.\end{matrix}$
We briefly examine Corollary 4.7 for these graphs. We have $I^{(2)}_{G}=I_{G}$ in (a) and (c), but not in (b). All generators of $I^{(2)}_{G}$ have degree $(1,\star)$, but $I_{G}$ of (b) has a generator of degree $(2,2)$.
We seek $\overline{\mathcal{V}}_{G}]$ for any copious graph $G$. The coefficients $\gamma_{i,G}$ can be found by solving polynomial equations as follows. We pick |G|i1|G|-i-1 random linear forms $\ell_{1}(s),\ldots,\ell_{|G|-i-1}(s)$ in the unknowns $s_{ij}$ for $ij\in G$. These define a linear subspace of dimension $i-\kappa$ in $\mathcal{K}_{G}$, where $\kappa=\kappa(G)$ from (7). Next, we pick $i-\kappa+2$ random linear forms $\ell^{\prime}_{1}(x),\ell^{\prime}_{2}(x),\ldots,\ell^{\prime}_{i-\kappa+2}(x)$ in the unknowns $x_{1},\ldots,x_{n}$. These forms define a linear subspace of dimension $\kappa+n-3-i$ in $\mathbb{P}^{n-1}$. The total number $|G|-\kappa+1$ of linear forms $\ell_{i}(s)$ and $\ell^{\prime}_{j}(x)$ equals the dimension of $\mathcal{V}_{G}$ in $\mathbb{P}^{|G|-1}\times\mathbb{P}^{n-1}$. As a result of this discussion, we have the following statement.
Lemma 4.10.
The multidegree of the scattering correspondence is given by the formula
$\gamma_{i,G}\,\,=\,\,\#\,\bigl\{\,(s,x)\in\mathcal{V}_{G}\,:\,\ell_{1}(s)=\,\cdots\,=\ell_{|G|-i-1}(s)\,=\,\ell^{\prime}_{1}(x)=\,\cdots\,=\ell^{\prime}_{i-\kappa+2}(x)\,=\,0\,\bigr\}.$
We now examine the case $n=6$. The $2^{15}$ subsets of $\binom{[6]}{2}$ yield $156$ non-isomorphic graphs. Most are not copious. For instance, only $17$ of the $156$ graphs satisfy the rank condition in (17). This includes the two graphs that were ruled out in Example 4.4. The following result summarizes our findings for $n=6$. The cases $n=7,8$ will be treated in Section 6.
The proof of Proposition 4.11 is by direct computation, mostly using Macaulay2. The column for $\overline{\mathcal{V}}_{G}]$ is found by the command multidegree. From this we get $\mu(G)$ by (27). The rightmost column lists the $\mathbb{Z}^{2}$-degrees of the minimal generators of the scattering ideal $I_{G}$. The index is the number of generators in that degree. For instance, the symbol $(12)_{3}$ means that the ideal $I_{G}$ has three minimal generators of degree $(1,2)$. The column $\delta(G)$ displays the degree of the logarithmic discriminant, which will be discussed in Section 8. The first graph in the list is the complete bipartite graph $K_{3,3}$. This is special because $\kappa(K_{3,3})=5$.
Proposition 4.11.
There are $15$ copious graphs for $n=6$. Their ideals $I_{G}$ are as follows:
[…]
For readers of [7] we note that $12$ of the $15$ copious graphs are not gentle. Their ideals $I_{G}$ have minimal generators of some degree $(2,\star)$ or $(3,\star)$. Here, the primality criterion in Corollary 4.7 does not apply, and it requires a non-trivial saturation step to compute $I_{G}$ from its subideal $I_{G}^{(2)}$. Adding to Example 4.4, here is another interesting non-copious graph.
Example 4.12.
[…]
$\begin{matrix}\quad x_{1}x_{2}x_{4}-x_{1}x_{2}x_{5}-x_{1}x_{3}x_{4}+x_{1}x_{3}x_{6}+x_{1}x_{4}x_{5}-x_{1}x_{4}x_{6}\\ +\,x_{2}x_{3}x_{5}-x_{2}x_{3}x_{6}-x_{2}x_{4}x_{5}+x_{2}x_{5}x_{6}+x_{3}x_{4}x_{6}-x_{3}x_{5}x_{6}.\,\end{matrix}$ (28)
Vanishing of this cubic is a necessary condition for a point in $\mathcal{M}_{0,6}$ to be a critical point of some scattering potential on $G$. We conclude that $\mathcal{V}_{G}=\emptyset$, meaning that $G$ is not copious.
One may wonder whether the class of copious graphs is closed under adding edges. The graph $K_{3,3}$ in the first row of Proposition 4.11 shows that this is not the case. Let $G^{\prime}$ be obtained from $K_{3,3}$ by adding one edge $ij$. The scattering matroid $\mathcal{S}(K_{3,3})$ has rank $8$, but $\mathcal{S}(G^{\prime})$ has rank $9$. The new edge $ij\in G^{\prime}$ is a co-loop, so condition (18) is violated. Namely, the kernel of each $S_{G}(u)$ is contained in a proper subspace of $\mathcal{K}_{G}$ given by $s_{ij}=0$. However, if we disallow bipartite graphs then the closure property holds. We state this without proof.
Corollary 4.13.
Let $G$ be a copious graph which is connected and not bipartite, and let $G^{\prime}$ be a graph obtained from $G$ by adding one edge. Then the larger graph $G^{\prime}$ is also copious.
5 Partial Compactifications
We shall derive a formula for $\mu(G)$ from Huh’s Theorem [6], which equates the signed Euler characteristic of a very affine variety with its ML degree. The very affine variety in our setting is a partial compactification $\mathcal{M}_{G}$ of the moduli space $\mathcal{M}_{0,n}$. We will now define this.
We follow the proof of Lemma 2.4. The kinematic lattice $\mathcal{K}_{G}\,\cap\,\mathbb{Z}^{|G|}$ is a free abelian group of rank $r=|G|-\kappa(G)$. Every vector $u\in\mathcal{K}_{G}\,\cap\,\mathbb{Z}^{|G|}$ gives rise to a character $\tilde{u}:(\mathbb{C}^{*})^{|G|}\rightarrow\mathbb{C}^{*}$. We write $\tilde{u}$ as a Laurent monomial in the $x_{i}-x_{j}$ for $ij\in G$. The generalized cross-ratio $\tilde{u}$ is constant along ${\rm PGL}(2)$ orbits on $\mathbb{C}^{n}$, so it is a regular function on the moduli space $\mathcal{M}_{0,n}$.
Let $u_{1},\ldots,u_{r}$ be a basis of $\mathcal{K}_{G}\cap\mathbb{Z}^{|G|}$. The generalized cross-ratios define a map
$\mathbb{P}^{n-1}\,\dashrightarrow\,(\mathbb{C}^{*})^{r},\,\,x\mapsto\bigl(\tilde{u}_{1}(x),\tilde{u}_{2}(x),\ldots,\tilde{u}_{r}(x)\bigr).$ (29)
This rational map descends to a well-defined regular map on the moduli space:
$\mathcal{M}_{0,n}\,\longrightarrow\,(\mathbb{C}^{*})^{r},\,\,x\mapsto\bigl(\tilde{u}_{1}(x),\tilde{u}_{2}(x),\ldots,\tilde{u}_{r}(x)\bigr).$ (30)
We define $\mathcal{M}_{G}$ as the Zariski closure in $(\mathbb{C}^{*})^{r}$ of the image of (29) or (30). Then $\mathcal{M}_{G}$ is a very affine variety. This means that $\mathcal{M}_{G}$ is the variety defined by an ideal $\mathcal{I}_{G}$ in the Laurent polynomial ring $\mathbb{C}[z_{1}^{\pm 1},z_{2}^{\pm 1},\ldots,z_{r}^{\pm 1}]$, whose elements are the regular functions on $(\mathbb{C}^{*})^{r}$. We note that $\mathcal{M}_{G}$ and $\mathcal{I}_{G}$ do not really depend on the choice of basis for the kinematic lattice $\mathcal{K}_{G}\,\cap\,\mathbb{Z}^{|G|}$. Passing to a different basis induces a change of coordinates on the torus $(\mathbb{C}^{*})^{r}$.
Example 5.1 (Octahedron).
Fix the graph $G$ in Example 2.5 and the basis (9) for its kinematic lattice. The map $\mathbb{P}^{5}\dashrightarrow(\mathbb{C}^{*})^{6}$ in (29) has the six coordinates $z_{i}=\tilde{u}_{i}(x)$ in (10). Elimination reveals that the very affine variety $\mathcal{M}_{G}$ is the complete intersection given by
$\mathcal{I}_{G}\,\,=\,\,\bigl\langle z_{1}z_{4}z_{5}-z_{4}-z_{5}+1,\,z_{2}z_{4}z_{6}-z_{2}z_{4}-z_{2}z_{6}+1,\,z_{3}z_{5}z_{6}-z_{3}z_{5}+z_{3}-z_{6}\bigr\rangle.$
Assuming momentum conservation (6), the scattering potential (11) can be written as
$L_{G}(t,z)\,=\,t_{1}\,{\rm log}(z_{1})\,+\,t_{2}\,{\rm log}(z_{2})\,+\,t_{3}\,{\rm log}(z_{3})\,+\,t_{4}\,{\rm log}(z_{4})\,+\,t_{5}\,{\rm log}(z_{5})\,+\,t_{6}\,{\rm log}(z_{6}),$
where the coefficients $t_{i}$ are linear combinations of the $s_{ij}$. We restrict this function to the subvariety $\mathcal{M}_{G}$ of $(\mathbb{C}^{*})^{6}$. Using Lagrange multipliers for $\mathcal{I}_{G}$, we find that this restriction has $\mu(G)=2$ critical points. Thus the threefold $\mathcal{M}_{G}$ has Euler characteristic $\chi(\mathcal{M}_{G})=-2$.
We now come to our fourth and last definition of what it means for a graph to be copious.
Definition 5.2.
A graph $G\subseteq\binom{[n]}{2}$ is topologically copious if the map (30) is an embedding. This property is independent of the choice of basis for the kinematic lattice $\mathcal{K}_{G}\,\cap\,\mathbb{Z}^{|G|}$.
Theorem 5.3.
For a topologically copious graph $G$, the very affine variety $\mathcal{M}_{G}$ is a partial compactification of $\mathcal{M}_{0,n}$, and the signed Euler characteristic of $\mathcal{M}_{G}$ is the ML degree $\mu(G)$.
Proof and Discussion.
The map (30) is an embedding of $\mathcal{M}_{0,n}$ into the torus $(\mathbb{C}^{*})^{r}$. The very affine variety $\mathcal{M}_{G}$ is the closure of $\mathcal{M}_{0,n}$ in $(\mathbb{C}^{*})^{r}$. This closure is a partial compactfication of $\mathcal{M}_{0,n}$. We discuss the structure of $\mathcal{M}_{G}$ in Section 7. For a similar construction see [5].
By Huh’s Theorem [6], the signed Euler characteristic of the very affine variety $\mathcal{M}_{G}$ equals the number of critical points of its log-likelihood function. That function is our scattering potential $L_{G}$, as seen from (8). Since $\mathcal{M}_{0,n}$ is open and dense in $\mathcal{M}_{G}$, for generic data $s\in\mathcal{K}_{G}$, all critical points of $L_{G}$ lie in $\mathcal{M}_{0,n}$. This proves the last assertion in Theorem 5.3. ∎
We now turn to a class of graphs $G$ for which $\mathcal{M}_{G}$ is the complement of a hyperplane arrangement. Equivalently, the very affine variety $\mathcal{M}_{G}\subset(\mathbb{C}^{*})^{r}$ is cut out by linear equations. A vertex of $G$ is a universal vertex if it is connected to all other vertices in $G$. This is the assumption made tacitly in many sources, including [4, 9, RS]. We state our result as follows.
Proposition 5.4.
Every copious graph that has a universal vertex is topologically copious.
Proof.
We write $ij]=x_{i}-x_{j}$ for the $2\times 2$ minors of our matrix which coordinatizes $\mathbb{P}^{n-1}$:
$X\,\,=\,\,\begin{pmatrix}1&1&1&\cdots&1\\ x_{1}&x_{2}&x_{3}&\cdots&x_{n}\end{pmatrix}$.
A projective isomorphism yields the coordinate system on $\mathcal{M}_{0,n}$ referred to as gauge fixing:
$Z\,\,=\,\,\begin{pmatrix}1&1&1&1&\cdots&1&0\\ 0&1&z_{3}&z_{4}&\cdots&z_{n-1}&1\end{pmatrix}$. (31)
The $2\times 2$ minors of $Z$ define an arrangement $\mathcal{A}$ of $\binom{n-1}{2}-1$ affine hyperplanes in $\mathbb{C}^{n-3}$. This is the coordinate choice used, e.g., in [4, equation (1)], [7, Section 3], and [15, equation (2)].
We can write the affine-linear forms for these hyperplanes as generalized cross-ratios:
$\frac{[1i][2n]}{[12][in]}=z_{i}\,,\quad\frac{[1n][2i]}{[12][in]}=z_{i}-1\,,\quad\frac{[1n][2n][ij]}{[12][in][jn]}\,=z_{j}-z_{i}\quad{\rm for}\,\,i,j\in\{3,4,\ldots,n-1\}.$ (32)
The corresponding elements of $\mathbb{Z}^{\binom{n}{2}}$ form a $\mathbb{Z}$-basis of the kinematic lattice $\mathcal{K}_{K_{n}}\cap\mathbb{Z}^{\binom{n}{2}}$:
$e_{1i}+e_{2n}-e_{12}-e_{in}\,,\,\,\,e_{1n}+e_{2i}-e_{12}-e_{in}\,,\,\,\,e_{1n}+e_{2n}+e_{ij}-e_{12}-e_{in}-e_{jn}.$ (33)
We now assume that $n$ is a universal vertex of $G$ and that $12$ is an edge of $G$. Our choice of $Z$ ensures that $[12]=1$ and $[in]=1$ for all $i$. Let $\mathcal{A}_{G}$ denote the subarrangement of $\mathcal{A}$ obtained by deleting all hyperplanes whose label $1i$, $2i$ or $ij$ is not an edge of $G$. The remaining linear forms in (33) form a $\mathbb{Z}$-basis of the kinematic space $\mathcal{K}_{G}$. The cross-ratios (32) for $ij\in G$ are regular functions on $\mathcal{M}_{G}$, written in the coordinates $z_{i}$ of $\mathbb{C}^{n-3}$.
We identified the very affine variety $\mathcal{M}_{G}$ with the arrangement complement $\mathbb{C}^{n-3}\backslash\mathcal{A}_{G}$. Since $G$ is copious, the map $\mathcal{V}_{G}\rightarrow\mathcal{K}_{G}$ is finite-to-one. By Varchenko’s Theorem, the ML degree $\mu(G)$ is the number of bounded regions of $\mathbb{R}^{n-3}\setminus\mathcal{A}_{G}$. In particular, $\mu(G)\geq 1$.
We fix one bounded region of $\mathbb{R}^{n-3}\setminus\mathcal{A}_{G}$, and we consider the affine-linear forms $z_{i}$, $z_{i}-1$ or $z_{j}-z_{i}$ which define the facets of that region. These span the space of all affine-linear forms on $\mathbb{R}^{n-3}$. In fact, since graphic matroids are unimodular, they contain a $\mathbb{Z}$-basis for the lattice of affine-linear forms on $\mathbb{Z}^{n-3}$. Hence, every coordinate $z_{3},z_{4},\ldots,z_{n-1}$ can be written as a polynomial in the generalized cross-ratios (32) that come from our $\mathbb{Z}$-basis of $\mathcal{K}_{G}\cap\mathbb{Z}^{|G|}$. This shows that the map $\mathcal{M}_{0,n}\rightarrow\mathcal{M}_{G}$ is injective, so $G$ is topologically copious. ∎
From the proof above, we can easily derive linear generators for the ideal $\mathcal{I}_{G}$ that defines $\mathcal{M}_{G}$ as a closed subvariety in $(\mathbb{C}^{*})^{r}$. Namely, we consider the $r$ expressions (32) which correspond to edges of $G$. These $r$ expressions are linear in the $z_{i}$. The relations among them generate $\mathcal{I}_{G}$, and these are linear as well. If $G$ does not have a universal vertex then $\mathcal{I}_{G}$ requires non-linear Laurent polynomials. We saw this for the octahedron in Example 5.1.
Remark 5.5.
[…]
We now conclude with the remaining step in the proof of our main theorem.
Proof of Theorem 1.3.
Suppose that $G$ is topologically copious, and consider the very affine variety $\mathcal{M}_{G}$ in $(\mathbb{C}^{*})^{r}$. The results in [6] ensure that the associated likelihood correspondence
$\mathbb{P}^{r-1}\,\longleftarrow\,\mathcal{V}_{G}\,\longrightarrow\,\mathcal{M}_{G}$
has the desired properties, namely the right map is dominant and the left map is $k$-to-11, where $k=|\chi(\mathcal{M}_{G})|\geq 1$. Since $\mathcal{M}_{0,n}$ is dense in $\mathcal{M}_{G}$, and $\mathcal{K}_{G}={\rm span}(u_{1},\ldots,u_{r})=\mathbb{P}^{r-1}$, the same holds for the maps in (5). This means that $G$ is geometrically copious. The other parts of Theorem 1.3 were established in Lemma 3.8, Lemma 4.5, and Proposition 5.4. ∎
6 Graphs with Few Vertices
The previous sections introduced four different ways for a graph $G$ to be copious. Theorem 1.3 tells us that they are equivalent. The only missing step is that copious implies topologically copious when $G$ has no universal vertex. We conjecture that this indeed holds. From now on, we call $G$ copious if it satisfies the equivalent conditions in Definitions 3.2, 3.7 and 4.2.
We undertook a comprehensive computational study of graphs with up to nine vertices. This section is dedicated to the findings of that study. The blueprint is given by the lists of copious graphs for $n=5$ in Example 4.9 and for $n=6$ in Proposition 4.11. We computed the analogous lists in the much harder cases $n=7,8$, with partial results also for $n=9$.
From the entry A000088 in the Online Encyclopedia of Integer Sequences (OEIS), we know that the number of graphs for $n=4,5,6,7,8,9$ equals $11,34,156,1044,12346,274668$. The data for the following theorem appears on our supplementary materials website [1].
Theorem 6.1.
The number of copious graphs for $n=4,5,6,7,8$ equals $1,3,15,129,2328$. All of these are topologically copious. For $n=9$, there are at least $26324$ copious graphs.
Proof and Discussion.
We obtain the lists of all graphs on $n=6,7,8,9$ vertices up to isomorphism from the graph library in SageMath. To identify the copious graphs, we first restrict to graphs that are compatible with momentum conservation. This means that the even cycle matroid $M(\mathcal{K}_{G})$ has no co-loop. We then compute the scattering matroid $\mathcal{S}(G)$ and check that it has no loops. Such a loop would prevent the map $\mathcal{V}_{G}\rightarrow\mathcal{K}_{G}$ from being dominant. At this stage we also check that the rank of $\mathcal{S}(G)$ satisfies inequality (20) up to $n=8$.
We wish to decide, for each remaining graph $G$, whether $G$ is matroidally copious. For this we need to check condition (18). We do this by a probabilistic method. Namely, we pick a reasonably large sample of points $u\in\mathbb{P}^{n-1}$ with rational coordinates. If $\kappa(G)$ equals the dimension on the left hand side of (18) for our sample, then we are done. Otherwise, we have strong evidence that $G$ is not copious. If so, it is best to try again with a larger sample.
To gain certainty, we turn to the algebraic definition of copious. We aim to compute the multidegree (26) for each graph $G$. We exclude $G$ whenever the ratio in (27) is not a positive integer. For $n=6$, we obtain all multidegrees symbolically with Macaulay2 within a few hours. For $n=7$, a significant fraction of the graphs require days. We therefore perform this task with numerical methods. Namely, we use the Julia package HomotopyContinuation.jl. This allows us to compute the multidegrees for $n=8$ as well.
The steps described in the first paragraph above yield $140$ graphs with loop-free scattering matroid for $n=7$. Among these, $11$ have a multidegree whose first coefficient is not divisible by $5$. This leaves us with $129$ graphs. To certify that these $129$ are matroidally copious, we checked (17) as well as (18). This was done in Macaulay2 with $21$ points $u$ sampled from $\mathbb{P}^{6}$. The same methodology also worked well for $n=8$ and $9$, giving a lower bound for the number of matroidally copious graphs. For $n=4,5,6,7,8$ we certified that the 1,3,15,129, 2328 graphs respectively in these lists are all the topologically copious and hence copious.
To verify that a graph is topologically copious, we again use Macaulay2. For a random point $x\in\mathcal{M}_{0,n},$ we check that the preimage of $(\tilde{u}_{1}(x),\tilde{u}_{2}(x),\dots,\tilde{u}_{r}(x))$ under the map (30) is a single PGL(2) orbit. This is done by computing an elimination ideal, but after fixing $x_{1}=1,\;x_{2}=2,\;x_{3}=3$. The resulting ideal for the preimage must then be a maximal ideal.
Our experimental results give us considerable confidence to conjecture that condition 4 is equivalent to conditions 1,2,3 in Theorem 1.3 also for graphs $G$ without a universal vertex.
Our method for identifying all matroidally copious graphs runs well also for larger $n$. But, it has the disadvantage that it is probabilistic. In practice, we can sample only finitely many $u\in\mathbb{P}^{n-1}$ when certifying the linear algebra conditions (18) and (19). By contrast, we found that computing multidegrees is infeasible for $n=9$ with the numerical method. ∎
We next present a more detailed view of the data in Theorem 6.1. Table 1 counts copious graphs $G$ for fixed $|G|$ and fixed $\mu(G)$. Each row and column corresponds to a range of values.
$\small\begin{matrix}\hskip 18.49988pt|G|\backslash\mu(G)\!\!\!\!&1&2&3&4&\!5,\!6\!&\!\!7..9\!\!&\!\!10..24\!\!\\ \!\!12&\!10&2&0&0&0&0&0\\ \!\!13&\!14&8&1&0&0&0&0\\ \!\!14&\!11&13&5&2&0&0&0\\ \!\!15&5&9&5&6&3&0&0\\ \!\!16&0&4&2&5&6&0&0\\ \!\!17&0&0&0&3&2&4&0\\ \!\!18..21&0&0&0&0&1&2&6\end{matrix}\qquad\,\,\,\,\footnotesize\begin{matrix}\qquad|G|\backslash\mu(G)\!\!\!\!&1&2&3,\!4&5,\!6&\!\!7..10\!\!&\!\!11..20\!\!&\!\!21..120\!\!\\ 13&0&1&0&0&0&0&0\\ 14&52&29&3&0&0&0&0\\ 15&\!109&95&39&1&0&0&0\\ 16&\!104&\!156&\!126&20&2&0&0\\ 17&52&\!133&\!185&89&15&0&0\\ 18&15&65&\!142&103&96&5&0\\ 19&0&16&63&63&114&50&0\\ 20,\!21&0&0&17&24&75&154&20\\ 22..28&0&0&0&0&0&23&72\end{matrix}\vskip-5.69046pt$
Table 1: Copious graphs by number of edges and ML degree for n=7n\!=\!7 (left) and n=8n\!=\!8 (right).
The first row for $n=8$ shows a unique graph $G$ with $|G|=13$. Interestingly, it has $\mu(G)=2$. This $G$ is the complete bipartite graph $K_{4,4}$ with three disjoint edges removed.
In the first column we see copious graphs $G$ with ML degree $\mu(G)=1$. If the edge set of $G$ is inclusion-maximal then this is a case of minimal kinematics [4]. The study in [4] was restricted to graphs with a universal vertex, with coordinates as in (31). Removing that universal vertex, the relevant graphs $H$ on [n1][n-1] are precisely the $2$-trees [4, Theorem 1.4].
According to A054581 in the OEIS, there are twelve $2$-trees for $n=8$. Each of them corresponds to a graph $G$ with 3n6=183n-6=18 edges. In Table 1 we see $15$ graphs with |G|=18|G|\!=\!18 and $\mu(G)\!=\!1$. Here is one of the three minimal kinematics graphs not coming from a $2$-tree.
Example 6.2.
Let $n=8$ and let $G$ be the graph with non-edges $12$, $13$, $14$, $18$, $23$, $24$, $27$, $34$, $36$, $45$. The graph $G$ does not contain a $2$-tree as a subgraph and its multidegree is
$6\sigma^{8}\tau^{5}+31\sigma^{9}\tau^{4}+52\sigma^{10}\tau^{3}+36\sigma^{11}\tau^{2}+10\sigma^{12}\tau+\sigma^{13}.$
Dividing the leading coefficient by n2=6n-2=6, equation (27) tells us that $\mu(G)=1$.
We state a conjecture which is consistent with our data and with the results in [4].
Conjecture 6.3.
Every minimal kinematics graph on $n]=\{1,2,\dots,n\}$ has 3n63n-6 edges.
Next comes a graph which is counted in the entry $2$ in the last row on the left in Table 1.
Example 6.4.
Let $n=7$ and consider the graph $G$ with non-edges $12,13,23$. This graph is copious with ${\rm rank}(\mathcal{S}(G))=11=7+(7-3)<18=|G|$. We describe the embedding (30) of $\mathcal{M}_{0,7}$ into $(\mathbb{C}^{*})^{11}$. We write $ij|k\ell]=\frac{(ik)(j\ell)}{(i\ell)(jk)}=\frac{(x_{i}-x_{k})(x_{j}-x_{\ell})}{(x_{i}-x_{\ell})(x_{j}-x_{k})}$ for the usual cross-ratio. For one choice of basis of the kinematic lattice $\mathcal{K}_{G}\cap\mathbb{Z}^{|G|}=\mathbb{Z}^{11}$, the embedding is given by
$\small\left([12|45],[12|46],[12|47],[13|45],[13|46],[13|47],[14|56],[14|57],[15|46],[15|47],\frac{[14][15][56]}{[16][17][45]}\right)\!.$
The multidegree of $\overline{\mathcal{V}}_{G}$ equals $40\sigma^{7}\tau^{4}+74\sigma^{8}\tau^{3}+45\sigma^{9}\tau^{2}+11\sigma^{10}\tau+\sigma^{11}$, and hence $\mu(G)=8$.
Our supplementary materials website [1] contains all graphs for $n=7,8,9$, together with our Julia and Macaulay2 code, as well as the datasets of ML degrees and multidegrees of scattering ideals. This is posted at the open-access repository Zenodo for research data:
https://zenodo.org/records/17546339.
7 Combinatorics of the ML degree
In this section, we discuss the number $\mu(G)$ of complex solutions to the graphical scattering equations (12) on a graph $G$. We conjecture that the following general formula holds.
Conjecture 7.1.
For any copious graph $G$ with $n$ vertices, the ML degree equals
$\mu(G)\;=\;\biggl|\,\sum_{s=3}^{n}(-1)^{\,s-3}\,\big(s-3\big)!\,\pi_{G}(s)\,\biggr|,$ (34)
where $\pi_{G}(s)$ is the number of partitions of the vertex set $[n]$ into $s$ independent subsets of $G$.
Using Theorem 6.1, we verified this conjecture for all graphs up to $n=8$ vertices.
Example 7.2 ($n=6$).
The $15$ graphs $G$ in Proposition 4.11 are presented by their non-edges. The number $\pi_{G}(s)$ counts set partitions of $\{1,2,3,4,5,6\}$ into $s$ blocks, where every pair in a block is a non-edge. The complete graph $K_{6}$ has $\pi_{G}(s)=0$ for $s\leq 5$ and $\pi_{G}(6)=1$. The octahedron graph, with non-edges $12$, $34$ and $56$, has $\pi_{G}(3)=\pi_{G}(6)=1$ and $\pi_{G}(4)=\pi_{G}(5)=3$, and the sum in (34) gives $\mu(G)=|1\cdot 1-1\cdot 3+2\cdot 3-6\cdot 1|=2$.
Example 7.3 (Two non-edges).
Let $G$ be the complete graph $K_{n}$ with two edges removed. If the two non-edges share a vertex, then (34) gives the formula $\mu(G)\,=\,(n-3)!-2(n-4)!$. But, if the two non-edges of $G$ are disjoint, then $\mu(G)\,=\,(n-3)!\,-\,2(n-4)!\,+\,(n-5)!$.
Example 7.4 (Removing a triangle).
Let $G$ be the graph on $[n]$ with non-edges $12,13,23$. Conjecture 7.1 states that $\mu(G)=(n-3)!-3(n-4)!+(n-5)!$. See Example 6.4 for $n=7$.
Discussion of Conjecture 7.1:.
We already conjectured that every copious graph is topologically copious. So, we assume the validity of that conjecture, and we consider graphs $G$ that are topologically copious. By Theorem 5.3, the ML degree $\mu(G)$ equals $(-1)^{n-3}$ times the Euler characteristic of the very affine variety $\mathcal{M}_{G}\subset(\mathbb{C}^{*})^{r}$. Hence Conjecture 7.1 is a topological statement, saying that the signed Euler characteristic of $\mathcal{M}_{G}$ is equal to (34).
The very affine variety $\mathcal{M}_{G}$ is a partial compactification of the moduli space $\mathcal{M}_{0,n}$. We think that $\mathcal{M}_{G}$ is obtained from $\mathcal{M}_{0,n}$ by adding a subset of strata $\mathcal{M}_{T}$ from the Deligne-Mumford compactification $\overline{\mathcal{M}}_{0,n}$. Each such stratum $\mathcal{M}_{T}$ is indexed by a phylogenetic tree $T$ whose leaves are labeled with $[n]$. If $T$ has $m$ interior nodes, and the $i$-th node has $s_{i}$ neighbors, then $\mathcal{M}_{T}$ is isomorphic to the product $\mathcal{M}_{0,s_{1}+1}\times\dots\times\mathcal{M}_{0,s_{m}+1}$. In particular, the codimension-$1$ strata are given by splits $\II\,\sqcup\,[n]\backslash I$ of $[n]$. If $|I|=k$ then this stratum is isomorphic to $\mathcal{M}_{0,k+1}\times\mathcal{M}_{0,n-k+1}$, so its Euler characteristic equals (k2)!(nk2)!(k-2)!(n-k-2)!.
We tried to identify all Deligne-Mumford strata in the boundary $\mathcal{M}_{G}\backslash\mathcal{M}_{0,n}$ and then add their Euler characteristics to (n3)!(n-3)!. Somehow this did not work out. The codimension-1 strata in the boundary are indexed by splits for which either $I$ or $\I[n]\backslash$ I is an independent set in $G$. The largest Euler characteristic among these arises when $I$ is a non-edge of $G$. This explains the subdominant term s=n1s=n-1 in the formula (34), and it proves the conjecture in the special case when the complement of $G$ is given by pairwise disjoint edges. ∎
Remark 7.5.
One approach to Conjecture 7.1 could be to study a more general question, which should be of independent interest. Consider any list of $r$ (generalized) cross-ratios, arising from lattice vectors $u_{1},\ldots,u_{r}$ in $\mathcal{K}_{K_{n}}$. This defines a regular map (30) from $\mathcal{M}_{0,n}$ to $(\mathbb{C}^{*})^{r}$. What can we say about this map? What are its fibers and its image? The closure of the image is a very affine variety. We can ask for its strata, its ML degree, and its prime ideal. If r=n3r=n-3 and $\tilde{u}_{1},\ldots,\tilde{u}_{r}$ are algebraically independent cross-ratios, then the map is dominant, and it is $\beta$-to-11 where $\beta$ is the cross-ratio degree studied by Silversmith [12].
Remark 7.6.
For graphs with a universal vertex as in Proposition 5.4, the partial compactification $\mathcal{M}_{G}$ coincides with the moduli space of graphically stable curves studied by Fry [5]. The boundary strata are parametrized by the dual graphs of such curves. Note that in [5] this moduli space is embedded into a torus via Plücker variables. Fry’s map coincides with our embedding for graphs with a universal vertex. See the proof of Proposition 5.4.
The formula in (34) is more general than the known formula for graphs with a universal vertex, i.e. for the complement of a graphic hyperplane arrangement, written as in (31). By Varchenko’s Theorem, the number of critical points equals the number of bounded regions. By Zaslavsky’s Theorem, this number can be determined from the characteristic polynomial of the underlying matroid. For graphic arrangements, this is the chromatic polynomial.
[…]
Theorem 7.7 (Varchenko-Zaslavsky).
[…]
Proof.
[…]
We now apply the relationship, found in [11, Proposition 3.1.2], between the characteristic polynomial of the deconed arrangement and that of the original graphical arrangement:
[…]
Example 7.8 ($n=6$).
Among the $15$ graphs in Proposition 4.11, only ten satisfy the hypothesis of Theorem 7.7. If $G$ is the graph with one non-edge $12$, then $G\backslash 6$ has nine edges, and its chromatic polynomial gives $\overline{p}_{G\backslash 6}(t)=t(t-2)(t-3)^{2}$. If $G$ has three non-edges $12,34,45$ then $\overline{p}_{G\backslash 6}(t)=t(t-2)(t^{2}-4t+5)$. The evaluations at $t=1$ are 4-4 and 2-2. Up to sign, these are the numbers seen in the column $\mu(G)$ for the table in Proposition 4.11.
8 Discriminants and Hypertrees
Our last section is devoted to two additional topics. The first is the logarithmic discriminant, which vanishes when $L_{G}$ has fewer than $\mu(G)$ critical points. This was studied by Kayser, Kretschmer and Telen [8] for the complete graph $K_{n}$ and it is here extended to any graph $G$. Second, we turn to hypertrees. This class of graphs was introduced by Castravet and Tevelev in [3] and its connection to scattering equations was featured in [4, Section 5]. We note that Questions 5.5 and 5.7 in [4] were our original motivation for launching this project. But, our investigations then took a different route, and these two questions remain largely open.
Let $G$ be a copious graph with $\mu(G)\geq 2$. The logarithmic discriminant $\Delta_{G}=\Delta_{G}(s)$ is an irreducible homogeneous polynomial of positive degree $\delta(G)$ in the unknowns $s_{ij}$. The hypersurface of $\Delta_{G}$ is the branch locus of the $\mu(G)$-to-11 map $(s,x)\mapsto s$ from the scattering correspondence $\mathcal{V}_{G}$ onto the kinematic space $\mathcal{K}_{G}$. Note that the logarithmic discriminant $\Delta_{G}$ is not unique as a polynomial in the $s_{ij}$. It is unique only modulo the linear relations (6).
We begin with $n=6$. The complete graph has $\mu(K_{6})=6$, and the degree of its discriminant is $\delta(K_{6})=30$. This was found numerically by Kayser, Kretschmer and Telen [8]. For details see Conjecture 1 and Example 8.5 in [8]. The value of $\delta(G)$ for other graphs $G$ is shown in our Proposition 4.11. We now focus on the octahedron (Examples 2.5 and 5.1).
The $12$ Mandelstam parameters $s_{ij}$ must satisfy momentum conversation at each of the six vertices, and there are many different choices of bases modulo these relations. There are four distinguished basis choices which respect the symmetry among the six vertices of $G$. Namely, we fix one of the four pairs of antipodal facets of the octahedron. Then the six edges not on those facets form a $6$-cycle, and these six $s_{ij}$ form a basis. Here is one of them:
$\small\begin{pmatrix}s_{14}\\ s_{16}\\ s_{46}\\ s_{23}\\ s_{25}\\ s_{35}\\ \end{pmatrix}\quad=\quad\frac{1}{2}\begin{pmatrix}-1&-1&-1&-1&\phantom{-}1&\phantom{-}1\,\,\\ -1&-1&\phantom{-}1&\phantom{-}1&-1&-1\,\,\\ \phantom{-}1&\phantom{-}1&-1&-1&-1&-1\,\,\\ -1&\phantom{-}1&\phantom{-}1&-1&-1&-1\,\,\\ \phantom{-}1&-1&-1&-1&-1&\phantom{-}1\,\,\\ -1&-1&-1&\phantom{-}1&\phantom{-}1&-1\,\,\end{pmatrix}\!\begin{pmatrix}s_{13}\\ s_{15}\\ s_{45}\\ s_{24}\\ s_{26}\\ s_{36}\end{pmatrix}$. (35)
These six linear equations are equivalent to (6). In the left column vector we see the antipodal facets $146$ and $235$, and the right column vector is the $6$-cycle 15426311-5-4-2-6-3-1.
To compute $\Delta_{G}$, we start with the equations $\partial L_{G}/\partial x_{i}=0$. Substituting (35), we obtain a system of bihomogeneous equations in $6$ unknowns $s_{ij}$ and $6$ unknowns $x_{i}$. Over a general point in $\mathcal{K}_{G}$, these equations define the union of two surfaces in $\mathbb{P}^{5}$ as in (14). By specializing three of the coordinates, say $x_{4}=1,x_{5}=2,x_{6}=3$, we obtain two points. We eliminate $x_{2},x_{3}$ from the system and obtain one quadratic equation in $x_{1}$ whose coefficients are expressions in the six parameters $s_{ij}$. The discriminant of this quadric in $x_{1}$ has the logarithmic discriminant $\Delta_{G}$ as its main factor. Extraneous factors are removed by changing the variable order in the elimination. Namely, $\Delta_{G}$ is the gcd of the discriminants of the three quadrics in $x_{1},x_{2},x_{3}$.
We find that the discriminant $\Delta_{G}$ is a homogeneous polynomial of degree six on $\mathcal{K}_{G}$:
$\Delta_{G}\,=\,s_{13}^{6}+2s_{13}^{5}s_{15}+2s_{13}^{5}s_{24}-2s_{13}^{5}s_{26}+2s_{13}^{5}s_{36}-2s_{13}^{5}s_{45}-s_{13}^{4}s_{15}^{2}+6s_{13}^{4}s_{15}s_{24}-6s_{13}^{4}s_{15}s_{26}+\,\cdots.$
This is a sum of $462$ monomials in the basis (35). This discriminant has remarkable properties. If we set one of the variables $s_{ij}$ to zero, then $\Delta_{G}(s)$ is the square of a cubic with $35$ terms. If we set two variables $s_{ij}$ to zero, then it becomes the product of three squares:
$\Delta_{G}|_{s_{13}=s_{45}=0}\,\,=\,\,(s_{15}-s_{24}+s_{26}+s_{36})^{2}\cdot(s_{15}-s_{24}-s_{26}-s_{36})^{2}\cdot(s_{15}+s_{24}+s_{26}-s_{36})^{2}.$
The discriminant $\Delta_{G}$ takes on both positive and negative values. For instance, $\Delta_{G}(s)=-71\,$ for $s_{13}=s_{15}=s_{24}=s_{26}=s_{45}=1,s_{36}=-2$. Here, both critical points of $L_{G}$ are non-real.
We found experimentally that the logarithmic discriminant $\Delta_{G}(s)$ of the octahedron $G$ is copositive when written in the $6$-cycle basis (35) for $\mathcal{K}_{G}$, i.e., $\Delta_{G}(s)>0$ for all $s\in\mathbb{R}_{>0}^{6}$. What makes this observation remarkable is that there is no gauge fixing for the octahedron. For all graphs $G$ with a universal vertex, copositivity follows immediately from [8, Corollary 4.2]. For instance, for $K_{5}$, this is the AM-GM inequality in [8, Example 1.1, page 3].
Problem 8.1.
Study the copositive geometry [14] of the logarithmic discriminant $\Delta_{G}(s)$ for all graphs $G$ that do not satisfy the hypothesis in Proposition 5.4 and Theorem 7.7.
We now discuss numerical computations of the discriminant degrees $\delta(G)$ for $n=6,7,8$. From [8] we already know that $\delta(K_{6})=30,\;\delta(K_{7})=208,\;\delta(K_{8})=1540$. For simplicity, we restrict our computations to those graphs that have a universal vertex, as in Proposition 5.4.
Proposition 8.2.
For the copious graphs $G$ on $[n]$ with a universal vertex, we found $\delta(G)\in\{4,10,16,30\}$ for $n=6$, $\delta(G)\in\{4,8,10,16,18,24,30,32,40,44,48,66,76,$ 84,104,144,208}84,104,144,208\} for $n=7$, and $4\leq\delta(G)\leq 1540$ for $n=8$, where $\delta(G)$ was always even.
Proof.
We use the Julia code due to Kayser, Kretschmer and Telen [8] for logarithmic discriminants. This is found at https://mathrepo.mis.mpg.de/LogDiscHyperplanes/. The computation for the complete graph $G=K_{n}$ is described in [8, Section 8]. It rests on the matrix $\mathcal{A}$ which encodes the braid arrangement, in the affine version of Remark 5.5.
For any graph $G$ with a universal vertex $n$, we remove the columns of the matrix $\mathcal{A}$ that are indexed by the non-edges of $G$. In some cases, we first performed a permutation of $[n]$ to ensure that the $n$ pairs $12,\,1n,2n,\ldots,(n-1)n$ are edges of $G$. This is necessary because [8] fixes the gauge $x_{1}=0,\;x_{2}=1,\;x_{n}=\infty$. Let $\tt s$} denote the resulting set of non-edges of $G$. This is a subset of $\binom{[n-1]}{2}$. We now run the following Julia code within the package of [8]:
      L = get_L_M0m(n)
      L’ = L[:, setdiff(1:size(L,2),s)]
      get_deg_DeltaLog(L’)
The output is the discriminant degree $\delta(G)$ of the copious graph $G$. The computation occasionally led to a numerical error, in which case we had to redo it more carefully. ∎
We now turn to hypertrees [4, Section 5]. We identify each hypertree $T$ with a graph $G$. Its edges are the pairs in the triples of $T$. The smallest such hypertree graph $G$ is the octahedron [4, eqn (26)], which belongs to a special family of hypertrees called bipyramids. The bipyramid graph $G$ has $n$ vertices and 3n63n-6 edges. Namely, we start with the (n2)(n-2)-cycle labelled by $3,4,5,\ldots,n,3$, and we connect both $1$ and $2$ with each vertex in the cycle. The following result can be viewed as a step towards answering [4, Question 5.5].
Proposition 8.3.
Every bipyramid is topologically copious, and hence copious.
Proof.
Let $G$ be the bipyramid on $[n]$ and write $\{e_{ij}\}_{1\leq i<j\leq n}$ for the standard basis of $\mathbb{Z}^{3n-6}$. Then $\{e_{13}+e_{2i}-e_{1i}-e_{23}\}_{i=4}^{n}$ is a linearly independent set in the kinematic space $\mathcal{K}_{G}$. We complete this to a basis $\{e_{13}+e_{2i}-e_{1i}-e_{23}\}_{i=4}^{n}\cup\{\tilde{u}_{n-2},\dots,\tilde{u}_{r}\}$ of the kinematic lattice $\mathcal{K}_{G}\cap\mathbb{Z}^{3n-6}$. Let $\Phi:\mathcal{M}_{0,n}\to(\mathbb{C}^{\ast})^{r}$ be the corresponding map (30). The generalized cross-ratios given by the basis elements $e_{13}+e_{2i}-e_{1i}-e_{23}$ are cross-ratios, and have the form
$\frac{(x_{1}-x_{3})(x_{2}-x_{i})}{(x_{1}-x_{i})(x_{2}-x_{3})}\,\,=\,\,[12|3i]\,=\,[21|i3].$
After fixing $x_{1}=\infty,x_{2}=0,x_{3}=1$, we find that $21|i3]=x_{i}$. Hence, the affine coordinates $x_{4},\ldots,x_{n}$ of $\mathcal{M}_{0,n}$ appear among the coordinates of $\Phi$. It follows that $\Phi$ is injective. ∎
Moving beyond bipyramids, this result holds for irreducible hypertrees with small $n$.
Proposition 8.4.
All irreducible hypertrees for $n\leq 10$ are copious.
Proof.
Using Macaulay2, we certified that all hypertrees in the Opie-Scheidwasser database https://websites.umass.edu/tevelev/hypertree_database/ are topologically copious. This was done with the method described in the proof of Theorem 6.1. ∎
Conjecture 8.5.
All irreducible hypertrees are copious.
We conclude our article with a simple formula for the ML degree of the bipyramid graph.
Corollary 8.6.
Assuming Conjecture 7.1, the ML degree of the bipyramid equals n4n-4.
Proof.
Any partition of $[n]$ into $s$ independent subsets either has $1$ and $2$ together in a part of size two or separately in two parts of size one. The number of such partitions is thus counted by partitions of the vertices of the (n2)(n-2)-cycle into s1s-1 and s2s-2 independent subsets. The chromatic polynomial of the $m$-cycle is $(t-1)^{m}+(-1)^{m}(t-1)$, and the number of partitions of a graph $H$ into $k$ independent sets equals $p_{H}(k)/k!$. This implies
$\pi_{G}(s)$ $\,\,=\,\,\,\frac{1}{(s-1)!}\sum_{i=0}^{s-1}(-1)^{s-1-i}\binom{s-1}{i}((i-1)^{n-2}+(-1)^{n-2}(i-1))$
$\quad+\,\,\frac{1}{(s-2)!}\sum_{j=0}^{s-2}(-1)^{s-j}\binom{s-2}{j}((j-1)^{n-2}+(-1)^{n-2}(j-1)).$
The ML degree $\mu(G)$ is the nested sum that arises by substituting $\pi_{G}(s)$ into (34). Our task is to show that this double sum is equal to n4n-4. A computer-aided proof for this identity was found by Manuel Kauers. We present this on our supplementary materials website [1]. ∎
Acknowledgement: We thank Manuel Kauers for his essential contribution to the proof of Corollary 8.6 and Benjamin Hollering for invaluable input on computations and for providing the list of graphs on $n=9$ vertices. We are grateful to Jake Levinson and Rob Silversmith for insightful discussions. BF and BZ were both supported by Graduate Research Fellowships in mathematics from the US National Science Foundation. VB and BS acknowledge support from the European Research Council (UNIVERSE PLUS, 101118787). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.